# Fundamental thresholds for computational and erasure errors via the coherent information

### v1

@@ III.2.2 Computational errors
We also consider errors occurring independently on each qubit given by the depolarizing channel:
@@
with $p=p_{x}+p_{y}+p_{z}$. The $n-$qubit error channel is then given by $\mathcal{N}=\prod_{i}\mathcal{N}_{i}$. Particularly we are interested in two choices of $(p_{x},p_{y},p_{z})$: i) Symmetric uncorrelated bit and phase flip noise: $p_{x}=p_{z}=p_{1}(1-p_{1})$ and $p_{y}=p_{1}^{2}$ with $p_{1}$ is the bit and phase flip probability; and ii) Symmetric depolarizing channel $p_{x}=p_{z}=p_{y}=p/3$. Unlike the erasure channel considered before, $\mathcal{N}$ is a CPTP map, therefore we must work with the mixed-state density matrices instead of a classical ensemble of states.
@@ III.3 Coherent information of QEC codes
where $S(\rho)=\operatorname{Tr}\rho\log\rho$ is the von Neumann entropy of the state $\rho$ and $\rho_{Q}=\operatorname{Tr}_{R}(\rho_{RQ})$. This quantity was first introduced as a measure of the amount of quantum information transmitted by a quantum channel Lloyd (1997) and as an indicator of the existence of a QEC protocol with maximum success probability Schumacher and Nielsen (1996). Recently, it has been shown that it can also signal optimal thresholds of QEC codes Fan et al. (2024); Wang et al. (2023); Hauser et al. (2024); Su et al. (2024); Lyons (2024); Zhao and Liu (2023); Colmenarez et al. (2024). The exact setting for studying optimal threhsolds of QEC codes is shown in Fig. 2. We start with a generalized Bell pair between the reference system $R$ and the code space of an error correcting code $[[n,k,d]]$ on the system $Q$ of $n$ qubits.
@@
where the state $|\bm{R}_{q}\rangle=|R_{1}R_{2}...R_{k}\rangle$ is a basis state of the reference system. The reference system is composed of $k$ qubits, each of them denoted by the quantum number $R_{i}$. A “code word” state $|\bm{L}_{q}\rangle=|L_{1}L_{2}...L_{k}\rangle$ is a basis $n-$qubit state for the code space of the QEC code living on the system $Q$, $L_{i}$ are the logical quantum numbers of the code space. The state $|\bm{S}\rangle=|S_{1},S_{2},...,S_{n-k}\rangle$ is the state for which all stabilizer quantum numbers (syndromes) are $S_{i}=+1$ $\forall i$. An alternative way to represent this state is
@@ IV Results
In this section, we describe in detail how to derive the mappings of the coherent information (CI) in the error configuration picture, both with and without qubit loss. In Sec. IV.1, we present the exact expressions for the CI of stabilizer codes with erasure errors. Then, in Sec. IV.2, we study computational errors by applying the methods used in Refs. Wang et al. (2023); Hauser et al. (2024); Fan et al. (2024); Lyons (2024) to CSS codes and depolarizing noise. In Sec. IV.3, we combine the methods and derivations explained in the two previous sections, and arrive at closed-form expressions for the CI of CSS codes under both computational and erasure errors.
@@ IV.1 Coherent information under erasure errors
In this section, we compute the CI of QEC codes under the action of the erasure channel given by Eq. 14. Recall that the erasure operation traces out $m=m(l)$ physical qubits in region $A_{l}$, where $l$ denotes the configuration of erased qubits. For a fixed erasure configuration $l$, we then obtain the state:
@@
Each erased qubit leads to a small set of stabilizers that must be redefined, while some others become ill-defined. In total, only $n-2m$ stabilizers remain well-defined after erasing $m$ qubits. The latter is a consequence of how projective measurements are treated in the stabilizer formalism Nielsen and Chuang (2010). For the surface/toric code, an example can be seen in Fig. 3. After erasure of one qubit, a total of four stabilizers is affected. Then each pair of $X$ and $Z$ stabilizers gets combined and becomes super-plaquette and super-star operators. Thus two out of four affected degrees of freedom remain well-defined. Furthermore, the fate of the logical operators also depends on the configuration $l$ of lost qubits. If at least one representative of the given logical operator still has support on the remaining physical qubits, then we say that such a logical operator remains well-defined. A logical operator becomes ill-defined once all representatives have support on at least one erased qubit. One example of the latter in the toric code is the case when erasure errors split the lattice into two disconnected parts, hence cutting the way of non-contractible loops, thereby impeding the existence of the respective logical (string) operators. To determine whether a logical operator $O_{L_{i}}$ remains well-defined, we use the procedure outlined in Ref. Amaro et al. (2020) (see Appendix A for details). If $h$ logical operators are not recoverable, then one can write
@@
$I=\sum_{l=1}P(l)I^{l}=k\log 2-\langle b_{l}\rangle\log 2-2\langle c_{l}\rangle\log 2.$ (26)
@@ IV.2.1 Calculation of $S(\rho_{RQ})$
$\rho_{RQ}^{r}=\sum_{C^{(s)}}\prod_{s=1}^{r}P(C^{(s)})w_{C^{(s)}}\rho^{0}_{RQ}w_{C^{(s)}},$ (28)
$\operatorname{Tr}(\rho_{RQ}^{r})=\sum_{C^{(s)}}\prod_{s=1}^{r}P(C^{(s)})\operatorname{Tr}\left(\prod_{s=1}^{r}w_{C^{(s)}}\rho^{0}_{RQ}w_{C^{(s)}}\right),$
$\operatorname{Tr}(\rho_{RQ}^{r})$ $=$ $\dfrac{1}{2^{kr}}\sum_{C^{(s)}}\sum_{q^{(s)},q^{{}^{\prime}(s)}}\prod_{s=1}^{r}P(C^{(s)})\operatorname{Tr}\left(\prod_{s=1}^{r}w_{C^{(s)}}|\bm{R}_{q^{(s)}},\bm{L}_{q^{(s)}},\bm{S}\rangle\langle\bm{R}_{q^{{}^{\prime}(s)}},\bm{L}_{q^{{}^{\prime}(s)}},\bm{S}|w_{C^{(s)}}\right).$ (30)
@@
$\operatorname{Tr}(\rho_{RQ}^{r})=\sum_{C^{(s)}}\prod_{s=1}^{r}P(C^{(s)})\langle\psi_{0}|w_{C^{(s)}}w_{C^{(s+1)}}|\psi_{0}\rangle,$ (31)
@@
$C^{(s+1)}=C^{(1)}+v^{(s)},\quad s=1,2,...,n-1.$ (32)
Here, $v^{(s)}$ stands for all error chains that do not flip any stabilizer quantum number $S_{i}$, i.e. a member of the stabilizer group. In the toric code they have an interpretation in terms of homologically trivial loops Dennis et al. (2002). One can further simplify this expression as
@@
$H[\{\eta^{x}\},\{\eta^{z}\}]=\sum_{\ell}\eta^{x}_{\ell}\left(J_{x}-\dfrac{J_{1}}{2}\right)P^{X}_{\ell}+$
$\eta^{z}_{\ell}\left(J_{z}-\dfrac{J_{1}}{2}\right)P^{Z}_{\ell}+\eta^{z}_{\ell}\eta^{x}_{\ell}J_{1}P^{Z}_{\ell}P^{X}_{\ell}.$ (43)
@@ IV.2.2 Calculation of $S(\rho_{Q})$
$\rho_{Q}^{r}=\sum_{C^{(s)}}\prod_{s=1}^{r}P(C^{(s)})w_{C^{(s)}}\rho^{0}_{Q}w_{C^{(s)}},$ (48)
$\operatorname{Tr}(\rho_{Q}^{r})=\sum_{C^{(s)}}\prod_{s=1}^{r}P(C^{(s)})\operatorname{Tr}\left(\prod_{s=1}^{r}w_{C^{(s)}}\rho^{0}_{Q}w_{C^{(s)}}\right),$ (49)
$\operatorname{Tr}(\rho_{Q}^{r})=\dfrac{1}{2^{kr}}\sum_{C^{(s)}}\sum_{\{q^{(s)}\}}\prod_{s=1}^{r}P(C^{(s)})\operatorname{Tr}\left(\prod_{s=1}^{r}w_{C^{(s)}}|\bm{L}_{q^{(s)}},\bm{S}\rangle\langle\bm{L}_{q^{(s)}},\bm{S}|w_{C^{(s)}}\right).$ (50)
@@
$S(\rho_{Q})$ $=$ $k\log 2+\log f-\dfrac{n}{4}\log\left[(1-p)p_{x}p_{y}p_{z}\right]$ (53)
@@
IV.2.3 Coherent information for depolarizing noise
@@
where $\mathcal{Z}_{0}$ is the partition function without defects along the support of logical operators. Importantly, Eq. (54) does not make any assumptions about the underlying details of the QEC code. Neither dependencies on $k$ nor the underlying geometry of the code are specified. Therefore, we expect that the same approach applies to LDPC codes or topological codes in any spatial dimension. We highlight that the details of the code enter via the variables $P^{X}_{\ell}$ and $P^{Z}_{\ell}$ defined in Eq. (37). Thus, the work lies in properly parameterizing these variables in terms of classical spins. In this work, we specifically tackle the 2D toric and color codes in Sec. V. For simplicity, we assume identical depolarizing channels on each qubit, but this can also be relaxed by allowing qubit-dependent couplings $J_{x},J_{z},J_{1}$ or considering correlated errors Chubb and Flammia (2021).
@@ IV.3.1 Calculation of $S(\rho_{RQ})$
$\rho_{RQ}^{r}=\sum_{C^{(s)}}\prod_{s=1}^{r}P(C^{(s)})w_{C^{(s)}}\rho^{l}_{RQ}w_{C^{(s)}},$ (56)
$\operatorname{Tr}(\rho_{RQ}^{r})=\sum_{C^{(s)}}\prod_{s=1}^{r}P(C^{(s)})\operatorname{Tr}\left(\prod_{s=1}^{r}w_{C^{(s)}}\rho^{l}_{RQ}w_{C^{(s)}}\right).$
@@
$\sum_{C^{(s)}}\prod_{s=1}^{r}P(C^{(s)})\langle\psi_{0}|w_{C^{(s)}}w_{C^{(s+1)}}|\psi_{0}\rangle,$
@@
$C^{(s+1)}=C^{(1)}+v^{(s)},\quad s=1,2,...,n-1.$ (59)
Here, $v^{(s)}$ represent all operations that do not change any of the well-defined $Z$ and $X$ stabilizer quantum numbers. In the case without erasures, all operators that commute with the $Z$ ($X$) stabilizers are members of the $X$ ($Z$) stabilizer subgroup. After erasing some qubits, the well-defined $X$ and $Z$ stabilizers cannot generate all error chains $C$. However, we must consider all $v^{(s)}$ that do not change the syndrome, even if they are not themselves made up of the well-defined $X$ and $Z$ stabilizers. Therefore, we can write
@@
The partition function $\mathcal{Z}_{l}\left[\{\eta^{x},\eta^{z}\}\right]=\sum_{\{P^{X(s)},P^{Z(s)}\}}\exp\left[H_{l}(s)\right]$ is the partition function of a classical Hamiltonian given by Eq. (IV.2.1) with the configuration $\{\eta^{x},\eta^{z}\}$ of couplings. Each choice of $\{\eta^{x},\eta^{z}\}$ represents an error chain, and $P(\{\eta^{x},\eta^{z}\})$ is its probability according to Table 2. The erasure errors enter as missing links with probability $e$ in the statistical mechanics model derived in Eq. IV.2. In general, erasure errors favor a fully symmetric equilibrium state by switching off the interaction on each link $\ell$ of the lattice. In Sec. V, we further discuss this model for 2D toric and color codes. Now, using the identity in Eq. (46), we obtain
@@ IV.3.2 Calculation of $S(\rho_{Q})$
$\operatorname{Tr}(\rho_{Q}^{r})$ $=$ $\frac{1}{2^{(k^{\prime}+b+m+c)r}}\sum_{C^{(s)}}\sum_{\{q^{(s)}\}}\sum_{\bm{S}^{{}^{\prime}(s)}}\sum_{\{o^{(s)}\}}\sum_{\{p^{(s)}\}}\prod_{s=1}^{r}P(C^{(s)})\operatorname{Tr}\Big{(}\prod_{s=1}^{r}w_{C^{(s)}}|\bm{L}_{q^{(s)}},\bm{S}^{(s)},\bm{S}^{{}^{\prime}(s)},\bm{L}_{o^{(s)}},\bm{L}_{p^{(s)}}\rangle$ (64)
@@
IV.3.3 Coherent information for depolarizing noise and erasure errors

### v2

@@ III.2.2 Computational errors
We also consider errors occurring independently on each qubit given by the Pauli channel:
@@
with $p=p_{x}+p_{y}+p_{z}$. The $n-$qubit error channel is then given by $\mathcal{N}=\prod_{i}\mathcal{N}_{i}$. Particularly we are interested in two choices of $(p_{x},p_{y},p_{z})$: i) Symmetric uncorrelated bit and phase flip noise: $p_{x}=p_{z}=p_{1}(1-p_{1})$ and $p_{y}=p_{1}^{2}$ with $p_{1}$ is the bit and phase flip probability; and ii) Symmetric depolarizing channel $p_{x}=p_{z}=p_{y}=p/3$.
@@ III.3 Coherent information of QEC codes
where $S(\rho)=-\operatorname{Tr}\rho\log\rho$ is the von Neumann entropy of the state $\rho$ and $\rho_{Q}=\operatorname{Tr}_{R}(\rho_{RQ})$. This quantity was first introduced as a measure of the amount of quantum information transmitted by a quantum channel Lloyd (1997) and as an indicator of the existence of a QEC protocol with maximum success probability Schumacher and Nielsen (1996). Recently, it has been shown that it can also signal optimal thresholds of QEC codes Fan et al. (2024); Wang et al. (2023); Hauser et al. (2024); Su et al. (2024); Lyons (2024); Zhao and Liu (2023); Colmenarez et al. (2024). The exact setting for studying optimal threhsolds of QEC codes is shown in Fig. 2. We start with a generalized Bell pair between the reference system $R$ and the code space of an error correcting code $[[n,k,d]]$ on the system $Q$ of $n$ qubits.
@@
where the state $|\bm{R}_{q}\rangle=|R_{1}R_{2}...R_{k}\rangle$ is a basis state of the reference system. The reference system is composed of $k$ qubits, each of them denoted by the quantum number $R_{i}$. A “code word” state $|\bm{L}_{q},\bm{S}\rangle=|L_{1}L_{2}...L_{k},S_{1},S_{2},...,S_{n-k}\rangle$ is a basis $n-$qubit state for the code space of the QEC code living on the system $Q$. $L_{i}$ are the logical quantum numbers of the code space and $S_{i}=+1$ $\forall i$ are the stabilizer quantum numbers (syndromes). An alternative way to represent this state is
@@ IV Results
In this section, we describe in detail how to derive the mappings of the coherent information (CI) in the error configuration picture, both with and without qubit loss. In Sec. IV.1, we present the exact expressions for the CI of stabilizer codes with erasure errors. Then, in Sec. IV.2, we study computational errors by applying the methods used in Refs. Wang et al. (2023); Hauser et al. (2024); Fan et al. (2024); Lyons (2024) to CSS codes and Pauli noise. In Sec. IV.3, we combine the methods and derivations explained in the two previous sections, and arrive at closed-form expressions for the CI of CSS codes under both computational and erasure errors. To simplify the notation across the section we choose to work with the base-two logarithm, such that $\log 2=1$.
@@ IV.1 Coherent information under erasure errors
In this section, we compute the CI of QEC codes under the action of the erasure channel given by Eq. 15. Recall that the erasure operation traces out $m=m(l)$ physical qubits in region $A_{l}$, where $l$ denotes the configuration of erased qubits. For a fixed erasure configuration $l$, we then obtain the state (the ancilla bit states $|a_{l}\rangle$ are omitted from now on):
@@
Each erased qubit leads to a small set of stabilizers that must be redefined, while some others become ill-defined. In total, only $n-k-2m$ stabilizers remain well-defined after erasing $m$ qubits. The latter is a consequence of how projective measurements are treated in the stabilizer formalism Nielsen and Chuang (2010). For the surface/toric code, an example can be seen in Fig. 3. After erasure of one qubit, a total of four stabilizers is affected. Then each pair of $X$ and $Z$ stabilizers gets combined and becomes super-plaquette and super-star operators. Thus two out of four affected degrees of freedom remain well-defined. Furthermore, the fate of the logical operators also depends on the configuration $l$ of lost qubits. If at least one representative of the given logical operator still has support on the remaining physical qubits, then we say that such a logical operator remains well-defined. A logical operator becomes ill-defined once all representatives have support on at least one erased qubit. One example of the latter in the toric code is the case when erasure errors split the lattice into two disconnected parts, hence cutting the way of non-contractible loops, thereby impeding the existence of the respective logical (string) operators. To determine whether a logical operator $O_{L_{i}}$ remains well-defined, we use the procedure outlined in Ref. Amaro et al. (2020) (see Appendix A for details). If $h$ logical operators are not recoverable, then one can write
@@
$I=\sum_{l=1}P(l)I^{l}=k-\langle b_{l}\rangle-2\langle c_{l}\rangle.$ (27)
@@ IV.2.1 Calculation of $S(\rho_{RQ})$
$\rho_{RQ}^{r}=\sum_{\{C^{(s)}\}}\prod_{s=1}^{r}P(C^{(s)})w_{C^{(s)}}\rho^{0}_{RQ}w_{C^{(s)}},$ (29)
$\operatorname{Tr}(\rho_{RQ}^{r})=\sum_{\{C^{(s)}\}}\Big{[}\prod_{s=1}^{r}P(C^{(s)})\Big{]}\operatorname{Tr}\left(\prod_{s=1}^{r}w_{C^{(s)}}\rho^{0}_{RQ}w_{C^{(s)}}\right),$
$\operatorname{Tr}(\rho_{RQ}^{r})$ $=$ $\dfrac{1}{2^{kr}}\sum_{\{C^{(s)}\}}\sum_{\{q^{(s)},q^{{}^{\prime}(s)}\}}\Big{[}\prod_{s=1}^{r}P(C^{(s)})\Big{]}\operatorname{Tr}\left(\prod_{s=1}^{r}w_{C^{(s)}}|\bm{R}_{q^{(s)}},\bm{L}_{q^{(s)}},\bm{S}\rangle\langle\bm{R}_{q^{{}^{\prime}(s)}},\bm{L}_{q^{{}^{\prime}(s)}},\bm{S}|w_{C^{(s)}}\right).$ (31)
@@
$\operatorname{Tr}(\rho_{RQ}^{r})=\sum_{\{C^{(s)}\}}\prod_{s=1}^{r-1}P(C^{(s)})\langle\psi_{0}|w_{C^{(s)}}w_{C^{(s+1)}}|\psi_{0}\rangle,$ (32)
@@
$C^{(s+1)}=C^{(1)}+v^{(s)},\quad s=1,2,...,r-1.$ (33)
Here, $v^{(s)}$ stands for all error chains that do not flip any stabilizer quantum number $S_{i}$ and commute with all logical operators, i.e. a member of the stabilizer group. In the toric code they have an interpretation in terms of homologically trivial loops Dennis et al. (2002). One can further simplify this expression as
@@
$H[\{\eta^{x}\},\{\eta^{z}\}]=\sum_{\ell}\eta^{x}_{\ell}\left(J_{x}-\dfrac{J_{1}}{2}\right)\dfrac{P^{X}_{\ell}}{2}+$
$\eta^{z}_{\ell}\left(J_{z}-\dfrac{J_{1}}{2}\right)\dfrac{P^{Z}_{\ell}}{2}+\eta^{z}_{\ell}\eta^{x}_{\ell}J_{1}\dfrac{P^{Z}_{\ell}P^{X}_{\ell}}{4}.$ (44)
@@ IV.2.2 Calculation of $S(\rho_{Q})$
$\rho_{Q}^{r}=\sum_{\{C^{(s)}\}}\prod_{s=1}^{r}P(C^{(s)})w_{C^{(s)}}\rho^{0}_{Q}w_{C^{(s)}},$ (49)
$\operatorname{Tr}(\rho_{Q}^{r})=\sum_{\{C^{(s)}\}}\Big{[}\prod_{s=1}^{r}P(C^{(s)})\Big{]}\operatorname{Tr}\left(\prod_{s=1}^{r}w_{C^{(s)}}\rho^{0}_{Q}w_{C^{(s)}}\right),$ (50)
$\operatorname{Tr}(\rho_{Q}^{r})=\dfrac{1}{2^{kr}}\sum_{\{C^{(s)}\}}\sum_{\{q^{(s)}\}}\Big{[}\prod_{s=1}^{r}P(C^{(s)})\Big{]}\operatorname{Tr}\left(\prod_{s=1}^{r}w_{C^{(s)}}|\bm{L}_{q^{(s)}},\bm{S}\rangle\langle\bm{L}_{q^{(s)}},\bm{S}|w_{C^{(s)}}\right).$ (51)
@@
$S(\rho_{Q})$ $=$ $k+\log f-\dfrac{n}{4}\log\left[(1-p)p_{x}p_{y}p_{z}\right]$ (54)
@@
IV.2.3 Coherent information for generic Pauli noise
@@
$I=k-\overline{\log\left[\dfrac{\sum_{D}\mathcal{Z}_{D}}{\mathcal{Z}_{0}}\right]}$ (55)
where $\mathcal{Z}_{0}$ is the partition function without defects along the support of logical operators. Importantly, Eq. (55) does not make any assumptions about the underlying details of the QEC code. Neither dependencies on $k$ nor the underlying geometry of the code are specified. Therefore, we expect that the same approach applies to LDPC codes or topological codes in any spatial dimension. We highlight that the details of the code enter via the variables $P^{X}_{\ell}$ and $P^{Z}_{\ell}$ defined in Eq. (38). Thus, the work lies in properly parameterizing these variables in terms of classical spins. In this work, we specifically tackle the 2D toric and color codes in Sec. V and the lift-connected surface code in Sec. VI. For simplicity, we assume identical Pauli channels on each qubit, but this can also be relaxed by allowing qubit-dependent couplings $J_{x},J_{z},J_{1}$ or considering correlated errors Chubb and Flammia (2021).
@@ IV.3.1 Calculation of $S(\rho_{RQ})$
$\rho_{RQ}^{r}=\sum_{\{C^{(s)}\}}\prod_{s=1}^{r}P(C^{(s)})w_{C^{(s)}}\rho^{l}_{RQ}w_{C^{(s)}},$ (57)
$\operatorname{Tr}(\rho_{RQ}^{r})=\sum_{\{C^{(s)}\}}\Big{[}\prod_{s=1}^{r}P(C^{(s)})\Big{]}\operatorname{Tr}\left(\prod_{s=1}^{r}w_{C^{(s)}}\rho^{l}_{RQ}w_{C^{(s)}}\right).$
@@
$\sum_{\{C^{(s)}\}}\prod_{s=1}^{r}P(C^{(s)})\langle\psi_{0}|w_{C^{(s)}}w_{C^{(s+1)}}|\psi_{0}\rangle,$
@@
$C^{(s+1)}=C^{(1)}+v^{(s)},\quad s=1,2,...,r-1.$ (60)
Here, $\{v^{(s)}\}$ represent all operations that do not change any of the well-defined $Z$ and $X$ stabilizer quantum numbers and commute with all logical operators. In the case without erasures, $v^{(s)}$ is always a member of the stabilizer group. After erasing some qubits, the well-defined $X$ and $Z$ stabilizers and logical operators cannot generate all error chains $C$. However, we must consider every $v^{(s)}$, even if they are not themselves made up of the well-defined $X$ and $Z$ stabilizers. Therefore, we can write
@@
The partition function $\mathcal{Z}_{l}\left[\{\eta^{x},\eta^{z}\}\right]=\sum_{\{P^{X(s)},P^{Z(s)}\}}\exp\left[H_{l}(s)\right]$ is the partition function of a classical Hamiltonian given by Eq. (IV.2.1) with the configuration $\{\eta^{x},\eta^{z}\}$ of couplings. Each choice of $\{\eta^{x},\eta^{z}\}$ represents an error chain, and $P(\{\eta^{x},\eta^{z}\})$ is its probability according to Table 2. The erasure errors enter as missing links with probability $e$ in the statistical mechanics model derived in Sec. IV.2. In general, erasure errors favor a fully symmetric equilibrium state by switching off the interaction on each link $\ell$ of the lattice. In Secs. V and VI, we further discuss this model for 2D toric and color codes and one class of LDPC code. Now, using the identity in Eq. (47), we obtain
@@ IV.3.2 Calculation of $S(\rho_{Q})$
$\operatorname{Tr}(\rho_{Q}^{r})$ $=$ $\frac{1}{2^{(k^{\prime}+b+m+c)r}}\sum_{\{C^{(s)}\}}\sum_{\{q^{(s)}\}}\sum_{\{\bm{S}^{{}^{\prime}(s)}\}}\sum_{\{o^{(s)}\}}\sum_{\{p^{(s)}\}}\Big{[}\prod_{s=1}^{r}P(C^{(s)})\Big{]}\operatorname{Tr}\Big{(}\prod_{s=1}^{r}w_{C^{(s)}}|\bm{L}_{q^{(s)}},\bm{S}^{(s)},\bm{S}^{{}^{\prime}(s)},\bm{L}_{o^{(s)}},\bm{L}_{p^{(s)}}\rangle$ (65)
@@
IV.3.3 Coherent information for Pauli noise and erasure errors